OpenStudy (anonymous):

Solve these system of equations. Please show your work. 2w - x + 5y + z = -3 3w + 2 x + 2y - 6z = -32 w + 3x + 3y - z =-47 5w -2 x - 3y + 3z = 49

OpenStudy (anonymous):

Heeey @ChmE sorry I had lost my internet connection

OpenStudy (anonymous):

Anyways...no I haven't studied matrices yet

OpenStudy (anonymous):

have you done matrices yet

OpenStudy (anonymous):

No i haven't :S

OpenStudy (anonymous):

ok so I'll teach the elimination method. This is the easiest when you have so many equations.

OpenStudy (anonymous):

Alright!

OpenStudy (anonymous):

There is a couple ways to do it. we can tackle to equations at a time, easier method. Or we can do all 4 at once, requires more thought but faster. What do you want to do

OpenStudy (anonymous):

two**

OpenStudy (anonymous):

I say we do two at a time for now.

OpenStudy (anonymous):

Lets take equations 1 and 3 to start. I chose these because they both have 1z and they happen to be opposites. We r going to add these equations together. 2w - x + 5y + z = -3 w + 3x + 3y - z = -47 3w + 2x + 8y = -50 (added 1 & 3)

OpenStudy (anonymous):

Are you still with me?

OpenStudy (anonymous):

Answers brought to you by http://www.bluebit.gr/matrix-calculator/linear_equations.aspx w=2 x=-12 y=-4 z=1 Call me back if you want to learn elimination

OpenStudy (anonymous):

Thanks!!